# UOE: Unlearning one Expert is Enough for Mixture-of-Experts LLMs

### v1

@@ 4 Our Proposal: Unlearning one Expert (UOE)
Uncovering the root cause: ‘short-cut’ in MoE LLM unlearning and expert selection shift. In order to fully understand the failure cases of MoE LLM unlearning, we begin by inspecting and monitoring the expert selection pattern of the unlearned model. In Fig. 2, we show the proportion of tokens assigned to each selected expert on the data samples from WMDP dataset (Li et al., 2024). For the input of a specific topic, a small portion of experts (around 6 to 9 out of 64 experts) were assigned with the majority of the tokens in each layer, which was also confirmed in Wang et al. (2024b). Thus, we have the following insight:
@@
(a) (b)
Figure 3: (a) Expert selection overlap ratio between the original pretrained model and the unlearned model with different unlearning iterations using GA on WMDP benchmark. (b) Forget loss vs. the number of unlearning iterations, when controlling parameters to unlearn in MoE LLM.
Based on the insight above, we define the frequently activated experts as topic-target experts, and the others as non-target. Thus, by eliminating the knowledge stored in these target experts, MoE LLM unlearning can be achieved more effectively. Next, we examine how the expert selection pattern evolves during unlearning. Specifically, we track the average expert selection overlap ratio across all layers between the unlearned model at different stages and the original pretrained model, when processing the forget set. The results, shown in Fig. 3 (a), reveal a steady decline in the overlap ratio as unlearning progresses, indicating that previously selected target experts are gradually replaced by non-target ones that do not contain the target knowledge. This shift persists even when routers are fixed, as unlearning can still indirectly influence router selection: a router’s decision at one layer depends on the output of the previous layer, which may have been affected by an updated expert of this previous layer in unlearning. Meantime, we observe a consistent reduction in forget loss, as shown in Fig. 3 (b). Thus, we can derive the following insight:
rgb]0.99,0.99,0.99 Insight 2: Existing unlearning methods tend to prompt the routers to shift expert selection from target to non-target experts unintentionally. This can create unlearning ‘shortcuts’ in expert selection to trick for low forget loss and lead to fake unlearning.
@@
Algorithm 1 UOE Unlearning Algorithm
@@
0:  Pretrained Model $\bm{\theta}_{o}$, Forget Set $\mathcal{D}_{f}$, Retain Set $\mathcal{D}_{r}$, Retain Loss $\ell_{r}$, Forget Loss $\ell_{f}$, Anchor Loss $L_{\text{anchor}}$
1:  $D_{s}\leftarrow$ Sample_Subset($\mathcal{D}_{f}$)
@@
3:  $s_{\text{top},l}\leftarrow$Ranking_And_Select($s$)
4:  Activate_Expert_And_Router($\bm{\theta}_{o},s_{\text{top}},$$\text{Router}^{l}$)
5:  $\bm{\theta}_{u}\leftarrow$Unlearn$(\bm{\theta}_{o},\ell_{f}(\mathcal{D}_{f}),\ell_{r}(\mathcal{D}_{r}),L_{\text{anchor}}^{(l)})$
@@
UOE for effective MoE LLM unlearning. As discussed earlier, a new paradigm tailored for MoE LLM unlearning is urgently needed to address the challenges of unintentional expert selection shifts in routers and excessive unlearning of non-target experts. Therefore, we propose a framework that (1) identifies the most relevant target experts, (2) ensures that these target experts remain highly activated throughout the unlearning process to avoid selection shifts, and (3) limits the impact of unlearning on non-target experts. Spurred by these, we introduce UOE, where unlearning is confined to a single expert. We refer the readers to Alg. 1 for an illustration of UOE. This approach starts with an expert attribution process to accurately identify the most relevant experts for the unlearning task.
Expert attribution. While the token assignment ratio for each expert, as shown in Fig. 2, can serve as a basic attribution metric, it overlooks finer details that are important for precise comparisons, due to the hidden states in each layer summed by weighted average. To address this, we adopt a gating score-based task affinity calculation method from (Wang et al., 2024b). Specifically, the affinity score for the $i$-th expert $e_{i}^{(l)}$ in the $l$-th layer of an MoE LLM is defined as:
$s_{i}^{(l)}=\frac{1}{Z}\sum_{j=1}^{Z}\frac{1}{L_{j}}\sum_{t=1}^{L_{j}}g_{i,t}^{(l)}$ (2)
where $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Z$ is size of the calibration dataset used for expert attribution, $L_{j}$ represents the length of the $j$-th input sequence $\mathbf{x}_{j}$, and $g_{i,t}^{(l)}$ is the probability score assigned to expert $\mathbf{e}_{i}^{(l)}$ for the $t$-th token. Following Wang et al. (2024b), the attribution data can be a subset universally sampled from the original forget set. We find that a subset containing over 100,000 tokens is robust enough to select the most relevant experts for an unlearning task. For each layer, we rank the experts based on their affinity score and select the top expert as the target expert for unlearning.
@@
$L_{\text{anchor}}^{(l)}=\|\mathbf{g}^{(l)}-[a_{1}^{(l)},a_{2}^{(l)},\dots,a^{(l)}_{E^{(l)}}]\|_{2}^{2},\,\,$ (3)
@@
where $E^{(l)}$ is the total number of experts in the $l$-th layer, $\mathbf{g}^{(l)}=[g^{(l)}_{1},g^{(l)}_{2},\dots,g^{(l)}_{i}]$ is the output of router, and $a_{i}^{(l)}=1$ if the $i$-th expert is identified as the target expert, otherwise $a_{i}^{(l)}=0$. The unlearning loss can be formularized as:
$\min_{\bm{\theta}}\ell_{f}(\bm{\theta};\mathcal{D}_{f})+\lambda\ell_{r}(\bm{\theta};\mathcal{D}_{r})+\alpha L_{\text{anchor}}^{(l)},$ (4)
@@
where $\alpha$ is a hyperparameter to control the strength of anchor loss. The sensitivity analysis of $\alpha$ is provided in Sec. D in Appendix.
A single-layer solution for MoE LLM unlearning by UOE. While we have successfully identified the most relevant target expert for each layer and implemented the router anchor loss to stabilize expert selection, applying unlearning across all layers still leads to expert selection shifts.
Figure 4: Expert selection overlap ratio for each MoE layer in DeepSeek-v2-Lite, comparing the unlearned and original models after 80 iterations with GA. The expert with the top affinity score in each layer is tunable during unlearning.
In Fig. 4, our results indicate that, even with the anchor loss, unlearning across multiple layers amplifies the effects of the MoE structure, where minor selection shifts in earlier layers are magnified, leading to substantial shifts in deeper layers. Consequently, the unlearned model still suffers a significant utility drop of over $30\%$ ($55.48\%$ before unlearning vs. $24.65\%$ after unlearning). To address this, we reduced the number of layers involved in the unlearning process. Surprisingly, unlearning just a single layer proved sufficient to achieve strong performance. By confining unlearning to one layer, we effectively minimize the cascading effects of expert selection shifts while still eliminating the target knowledge. In practice, we select the layer with the highest top-1 expert affinity score, as calculated in (2), and perform unlearning on its corresponding expert, which is treated as the target expert for the entire model. Since LLM architectures are consistent in size across layers, the highest affinity scores per layer can be directly compared. This approach forms the foundation of UOE, allowing it to maintain model utility and forget effectiveness with exceptional efficiency. To validate the effectiveness of the current design, in Sec. 5, we conduct extensive discussions and empirical studies on various other possible design choices than UOE. These include unlearning multiple experts in a single layer, unlearning single experts across multiple layers, and exploring different expert selection schemes beyond the affinity-score-based approach.
5 Experiment
5.1 Experiment Setups
Unlearning tasks and datasets. To demonstrate the effectiveness of our proposed method, we evaluate and compare it against different baselines on two widely accepted LLM unlearning benchmarks: WMDP (Li et al., 2024) and RWKU (Jin et al., 2024). WMDP assesses the model’s ability to unlearn and prevent the generation of hazardous knowledge in biosecurity, cybersecurity, and chemical security contexts. RWKU, on the other hand, evaluates the model’s capability to eliminate knowledge about 200 real-world celebrities, simulating a private information protection task. We note that other commonly used benchmarks, such as TOFU (Maini et al., 2024) and MUSE (Shi et al., 2024), are less appropriate in this work. These benchmarks require models to be fine-tuned before unlearning, which introduces additional biases to the results for MoE LLMs due to the known instability in training and the tricky hyper-parameter tuning involved (Jiang et al., 2024), often leading to training collapse (Zoph et al., 2022a).
Models. We evaluate different unlearning methods on two MoE LLMs: Qwen1.5-MoE-A2.7B-Chat (Qwen), mistralai/Mixtral-8x7B-Instruct-v0.1 (Mixtral), and DeepSeek-V2-Lite (DeepSeek), representing the two mainstream MoE LLM training schemes: upcycle-from-dense and train-from-scratch, respectively. Qwen has a total of 14.3 billion parameters, with 2.7 billion activated during inference, while DeepSeek has 16 billion parameters, of which 2.4 billion are activated during inference. Mixtral has 45 billion parameters, of which 12.9 billion are activated. Due to the computation limitation, Mixtral is only applied on parameter-efficient fine-tunin settings.
Evaluation setup. We evaluate the performance of the unlearned LLMs based on two key metrics: unlearning efficacy (UE) and preserved model utility (UT). For the WMDP task, UE is measured using the WMDP-Cyber subsets provided by the benchmark. Specifically, we use forget accuracy (FA)—the accuracy of the LLMs on the forget set after unlearning—as the measure of UE. A lower FA indicates better unlearning. Given the four-option multiple-choice format of the test set, the ideal FA is $0.25$, equivalent to random guessing. UT is assessed using the zero-shot accuracy on the MMLU dataset (Hendrycks et al., 2020), which reflects the model’s ability to retain general knowledge. For the RWKU task, we use the Rouge-L recall score to evaluate performance on fill-in-the-blank and question-answer tasks, with lower scores indicating more effective unlearning. Since the task follows a question-answer format, the ideal FA is $0.0$, indicating no overlap between the generated answer and the ground truth. The UT evaluation for RWKU is the same as for WMDP, using the MMLU benchmark. By default, during the unlearning process, we select the model checkpoint that achieves the best balance between UE and UT as the optimal checkpoint.
Baselines. We demonstrate the effectiveness of our proposed UOE framework by comparing it against the LLM unlearning baselines: Gradient Ascent (GA(Eldan & Russinovich, 2023), Gradient Difference (GDiff(Maini et al., 2024) and most recent unlearning algorithm Negative Preference Optimization (NPO(Zhang et al., 2024) and Representation Misdirection for Unlearning (RMU(Li et al., 2024). For each method, we compare the original results with those obtained when incorporating UOE. Given the parameter efficiency of UOE, we also compare it with two state-of-the-art parameter-efficient fine-tuning (PEFT) methods for MoE LLMs: the low-rank adaptation scheme (LoRA) (Hu et al., 2021) and the Expert-Specialized Fine-Tuning method (ESFT) Wang et al. (2024b), which is specifically designed for MoE LLMs.
5.2 Experiment Results
Table 2: Performance comparison of existing unlearning methods equipped w/ and w/o UOE in WMDP (Li et al., 2024) and RWKU Jin et al. (2024) benchmarks on two MoE LLMs, namely Qwen1.5-MoE-A2.7B-Chat (Qwen) Team (2024) and DeepSeek-V2-Lite (DeepSeek) (Dai et al., 2024). The $\uparrow$ and $\downarrow$ symbols denote metrics where higher/lower values are better. The occurrence of significant utility drop (over $10\%$ drop in UT compared to the pretrained model) are marked in red.
@@ 5.2 Experiment Results
RMU+UOE $0.2536$ $0.5351$ $0.2859$ $0.5424$ $0.0723$ $0.5975$ $0.0130$ $0.5388$
Effectiveness of UOE in preserving model utility and unlearning efficacy. In Tab. 2, we present the UE (unlearning efficacy) and UT (utility) performance of our proposed UOE when integrated into different unlearning methods GA, GDiff, NPO, and RMU. First, one of the most notable findings is that UOE significantly improves model utility (UT) across all tested methods. For instance, when applied to baseline methods like GA, GDiff, and RMU, UOE consistently mitigates the severe utility drops (greater than $10\%$) that occur with the unmodified methods. This is particularly evident in scenarios where baseline methods without UOE exhibit drastic performance degradation in model utility (highlighted in red), while the same methods paired with UOE show substantial recovery. For example, the utility of GA on Qwen for the WMDP task drops from $0.5979$ to $0.3393$, but with UOE, the utility improves to $0.5012$, restoring much of the lost performance. Similarly, GDiff on RWKU suffers a significant utility loss from $0.5979$ to $0.3495$, but when UOE is applied, utility rises back to $0.5253$, nearly matching the original pretrained performance. Second, beyond utility preservation, the unlearning efficacy (UE)—measured through FA—remains either unaffected or slightly improved when UOE is employed. This balance between utility preservation and effective unlearning highlights the advantage of UOE. For instance, GDiff+UOE reduces the forget accuracy (FA) on Qwen (WMDP) from $0.2964$ to $0.2445$, demonstrating better unlearning while still achieving a higher utility score. Similarly, RMU+UOE on DeepSeek (WMDP) lowers the FA from $0.2530$ to $0.2859$, with a corresponding utility improvement from $0.4540$ to $0.5424$. Notably, methods such as GDiff and RMU, which experience significant utility loss when used alone, benefit greatly from the application of UOE, achieving near-pretrained utility levels while still maintaining effective unlearning.
Table 3: Tunable parameter ratio, PEFT vs UOE.

### v2

@@ 4 Our Proposal: SEUF
Figure 3: (left) Overlap ratio of selected experts between the original pretrained model and the unlearned model with different unlearning iterations using GA on WMDP benchmark. (right) Forget loss vs. the number of unlearning iterations, when controlling parameters to unlearn in MoE LLM.
Based on the insight above, we define the frequently activated experts as topic-target experts, and the others as non-target. Thus, by eliminating the knowledge stored in these target experts, MoE LLM unlearning can be achieved more effectively.
Next, we examine how the expert selection pattern evolves during unlearning. Specifically, we track the average expert selection overlap ratio across all layers between the unlearned model at different stages and the original pretrained model, when processing the forget set. The results, shown in Fig. 3 (a), reveal a steady decline in the overlap ratio as unlearning progresses, indicating that previously selected target experts are gradually replaced by non-target ones that do not contain the target knowledge. This shift persists even when routers are fixed, as unlearning can still indirectly influence router selection: a router’s decision at one layer depends on the output of the previous layer, which may have been affected by an updated expert of this previous layer in unlearning. Meantime, we observe a consistent reduction in forget loss, as shown in Fig. 3 (b). Thus, we can derive the following insight:
rgb]0.99,0.99,0.99 Insight 2: Existing unlearning methods tend to prompt routers to shift selection from target to non-target experts unintentionally. This creates unlearning ‘shortcuts’ in expert selection to trick for low forget loss and lead to fake unlearning.
@@
SEUF for effective MoE LLM unlearning. As discussed earlier, a new paradigm tailored for MoE LLM unlearning is urgently needed to address the challenges of unintentional expert selection shifts in routers and excessive unlearning of non-target experts. Therefore, we propose a framework that (1) identifies the most relevant target experts, (2) ensures that these target experts remain highly activated throughout the unlearning process to avoid selection shifts, and (3) limits the impact of unlearning on non-target experts. Spurred by these, we introduce SEUF, where unlearning is confined to $M$ most relevant target experts. We refer the readers to Alg. 1 for an illustration of SEUF.
This approach starts with an expert attribution process to accurately identify the most $M$ relevant experts for the unlearning task (step 1-3). Then, the gradient computation selected experts $e_{M}$ and their corresponding routers $\text{R}_{e_{M}}$ are enabled (step 4), while other parameters are frozen. Step 5 performs unlearning using any unlearning approach, as our framework is flexible. For example, gradient ascent can be applied with our defined loss functions. Next, we present the details of the expert attribution process and define the anchor loss function.
Algorithm 1 SEUF Unlearning Algorithm
@@
0:  Pretrained model $\bm{\theta}_{o}$, forget set $\mathcal{D}_{f}$, retain set $\mathcal{D}_{r}$, Setup: Retain loss $\ell_{r}$, forget loss $\ell_{f}$, anchor loss $L_{\text{anchor}}$, the number of experts to select $M$
1:  $\mathcal{D}_{s}\leftarrow$ Sample_Subset($\mathcal{D}_{f}$)
@@
3:  $e_{M}\leftarrow$Ranking_And_Select($s,M$)
4:  Activate_Expert_And_Router($\bm{\theta}_{o},e_{M},$$\text{R}_{e_{M}}$)
5:  $\bm{\theta}_{u}\leftarrow$Unlearn$(\bm{\theta}_{o},\ell_{f}(\mathcal{D}_{f}),\ell_{r}(\mathcal{D}_{r}),L_{\text{anchor}})$
@@
Expert attribution. While the token assignment ratio for each expert (shown in Fig. 2), can serve as a basic attribution metric, it overlooks finer details that are important for precise comparisons, due to the hidden states in each layer summed by weighted average. To address this, we adopt a gating score-based task affinity calculation method from (Wang et al., 2024b). Specifically, the affinity score for the $i$-th expert $e_{i}^{(l)}$ in the $l$-th layer of an MoE LLM is defined as:
$\vspace*{-0.05in}s_{i}^{(l)}=\frac{1}{Z}\sum_{j=1}^{Z}\frac{1}{L_{j}}\sum_{t=1}^{L_{j}}g_{i,t}^{(l)}\vspace*{-0.05in}$ (2)
where $Z$ is size of the calibration dataset used for expert attribution, $L_{j}$ represents the length of the $j$-th input sequence $\mathbf{x}_{j}$, and $g_{i,t}^{(l)}$ is the probability score assigned to expert $\mathbf{e}_{i}^{(l)}$ for the $t$-th token. Following Wang et al. (2024b), the attribution data can be a subset universally sampled from the original forget set. We find that a subset containing over 100,000 tokens is robust enough to select the most relevant experts for an unlearning task. For each layer, we rank the experts based on their affinity score and then finally select the top $M$ experts as the target expert for unlearning ($e_{M}$ in Algo. 1).
@@
$\vspace{-0.05in}L_{\text{anchor}}^{(l)}=\|\mathbf{g}^{(l)}-[a_{1}^{(l)},a_{2}^{(l)},\dots,a^{(l)}_{E^{(l)}}]\|_{2}^{2},\,\,$ (3)
@@
where $E^{(l)}$ is the total number of experts in the $l$-th layer, $\mathbf{g}^{(l)}=[g^{(l)}_{1},g^{(l)}_{2},\dots,g^{(l)}_{i}]$ is the output of router, and $a_{i}^{(l)}=1$ if the $i$-th expert is identified as the target expert, otherwise $a_{i}^{(l)}=0$. The unlearning loss can then be formularized as:
$\vspace*{-0.05in}\min_{\bm{\theta}}\ell_{f}(\bm{\theta};\mathcal{D}_{f})+\lambda\ell_{r}(\bm{\theta};\mathcal{D}_{r})+\alpha L_{\text{anchor}}^{(l)},\vspace*{-0.05in}$ (4)
@@
where $\alpha$ controls the strength of anchor loss. Its sensitivity is analyzed in Appendix Sec. B.
Selection of top $M$ experts. When forming $e_{M}$ of the top $M$ experts, there are two approaches: 1) selecting the top $M$ experts from all experts across all layers based on the affinity score $s_{i}^{(l)}$ in Eq.2; and 2) to mitigate selection shift from previous layers, another approach is to choose the top $M$ experts from the same layer. We examined both approaches under different settings $M$=1,3,6, and present the results in Tab. 2. We observe that unlearning a single expert ($M$=1) yields better performance than unlearning multiple experts, regardless of whether they come from the same layer or different layers. This trend of single-expert unlearning yielding the best performance is also observed across other unlearning tasks (see Tab. 7 in Appendix). This suggests:
rgb]0.99,0.99,0.99 Insight 4: Unlearning top-1 expert is the most effective.
[…]
Selected experts Top-1 Top-3 Top-6 Top-1+4-shared
Same layer 0.5100 0.4856 0.4652 0.3554
Different layers 0.2852 0.2567
From Tab. 2, we also observe that unlearning multiple experts across different layers leads to a substantial performance decline. To further analyze the Insight 4, let the total gradient update during unlearning be: $\Delta W=\sum_{i\in e_{M}}\lambda_{i}\nabla\mathcal{L}_{i},$ where $e_{M}$ is the set of selected experts being unlearned, $\lambda_{i}$ denotes their contribution weight, and $\nabla\mathcal{L}_{i}$ is their corresponding gradient update in Eq. (4). When only the top-1 expert is selected for unlearning, the modification to the weights remains minimal, ensuring low gradient interference. For multiple experts within the same layer, the gradient updates may partially cancel out, leading to moderate disruption. However, for multiple experts across different layers, the gradient updates affect distinct feature hierarchies, resulting in an unstable gradient flow and widespread model disruption.
This analysis also explains the deficiency of unlearning shared experts. In a given layer, shared experts are activated for all tokens, making them intuitively suitable targets for unlearning. However, Tab. 2 shows that unlearning the top-1 expert along with 4 shared experts causes a greater utility drop than unlearning top-6 experts in the same layer. Shared experts influence a broader range of token representations, so making them active for unlearning triggers high-magnitude gradient updates across multiple pathways. Also, since shared experts consolidate common knowledge across diverse contexts Liu et al. (2024a), their modification disrupts the model more severely, making them suboptimal for unlearning.
Table 3: Performance comparison of existing unlearning methods equipped w/ and w/o SEUF on WMDP (Li et al., 2024) and RWKU Jin et al. (2024) benchmarks on two MoE LLMs, namely Qwen1.5-MoE-A2.7B-Chat (Qwen) Team (2024) and DeepSeek-V2-Lite (DeepSeek) (Dai et al., 2024). Additionally, a group of baselines applying PEFT (LoRA and ESFT) on GA is included to evaluate our method’s effectiveness in selecting a suitable subset of parameters for unlearning, along with a baseline using random expert selection with RMU. The occurrence of significant utility increase (over $5\%$ increase in UT compared to without SEUF) are marked in green.
@@
RMU+SEUF $0.2536$ $0.5351$ $0.2859$ $0.5424$ $0.0723$ $0.5975$ $0.0130$ $0.5388$
GA+LoRA 0.2459 0.2689 0.2657 0.2295 $0.0000$ $0.2689$ 0.0000 0.2302
GA+ESFT 0.3145 0.4514 0.2737 0.5108 0.001 0.4433 0.0200 0.5001
RMU+Random 0.3505 0.5947 0.2722 0.5183 0.2110 0.5924 0.1176 0.5182
5 Evaluation Experiments
To demonstrate the effectiveness of our proposed method, we evaluate and compare it against different baselines on two widely accepted LLM unlearning benchmarks: WMDP (Li et al., 2024) and RWKU (Jin et al., 2024). The detailed experimental setup, such as unlearning tasks, datasets selection, targeted MoE models, unlearning baselines and hyper-parameter setting, is provided in Appendix Sec. A, due to space limitation. We next present results of several key experiments.
Effectiveness of SEUF across benchmarks and unlearning methods. In Tab. 3, we present the FE (forget efficacy) and UT (utility) of our proposed SEUF when integrating different unlearning methods GA, GDiff, NPO, and RMU. In this evaluation, SEUF selects only the top-1 expert for unlearning. There are two notable findings. First, SEUF effectively enhances unlearning, either by further reducing FE or maintaining a similar level compared to baselines without SEUF. Second, SEUF consistently improves model utility (UT) across all tested methods. Notably, for methods where UT drops by more than 10% (compared to the pretrained model), highlighted in red, SEUF mitigates the decline. For example, the utility of GA on Qwen for the WMDP task drops from $0.5979$ to $0.3393$, but with SEUF, the utility improves to $0.5012$, This demonstrates SEUF’s effectiveness in balancing unlearning performance and model retention. Notably, methods such as GDiff and RMU, which experience notable utility loss when used alone, benefit greatly from the application of SEUF, achieving near-pretrained utility levels while still maintaining effective unlearning.
SEUF outperforms parameter-efficient fine-tuning (PEFT) methods when used for unlearning. Tab.  3 also includes a set of baselines that apply PEFT on GA. It is used to evaluate whether our method unlearns more effectively a subset of parameters (top-1 expert) compared to PEFT. Tab. 4 shows a comparison of the parameter efficiency involved in tuning. The key conclusion from these results is: SEUF achieves far better parameter efficiency, with only $0.06\%$ of tunable parameters, compared to LoRA ($0.92\%$) and ESFT ($2.86\%$), while still maintaining a comparable level of forget efficacy and outperforming them in utility preservation. For instance, in RWKU, GA+SEUF achieves utility scores of 0.5709 on Qwen and 0.5485 on DeepSeek, significantly higher than LoRA (0.2689 and 0.2302) and ESFT (0.4433 and 0.5001).
Table 4: Tunable parameter ratio, PEFT vs SEUF.